OpenStudy (anonymous):

Problem . Solve the semi-infinite plate problem if the bottom edge of width \(\ \pi\) is held at T=\(\\cos x \), and the other sides are at 0o

OpenStudy (anonymous):

Solve the semi-infinite plate problem if the bottom edge of width \(\ \pi\) is held at T=\(\\cos x \), and the other sides are at 0o

OpenStudy (anonymous):

this is the same problem gerryliyana, good practice for u :)

OpenStudy (anonymous):

ok.., i'll try :)

OpenStudy (anonymous):

the only thing has been changed is boundary condition of bottom edge and length of wall

OpenStudy (anonymous):

Ok.., ok, i've tried. and i got \[T (x,0) \sum_{n=1}^{\infty} a_{n} \sin (nx)\] then into fourier form: \[a_{n} \frac{ 2 }{ l } \int\limits_{0}^{l} f(x,0) \sin (nx) dx\] because of T = f(x,0)= cos x for y =0; then \[a_{n} = \frac{ 2 }{ \pi } \int\limits_{0}^{\pi} \cos x . \sin (nx) dx\] how to solve \(a_{n} = \frac{ 2 }{ \pi } \int\limits_{0}^{\pi} \cos x . \sin (nx) dx\) ?? Would you kinly help me guys ??

OpenStudy (anonymous):

@ajprincess @.Sam. @ganeshie8 @jim_thompson5910 would you kinly help me ??

OpenStudy (anonymous):

@kropot72